What is the equation of a line perpendicular to 4x+2y=24?
Answer by alanc(27)

given line : 4x + 2y = 24. Convert it to Slope intercept form.
2y = -4x + 24
y = -2x + 12
Here our slope is -2
a line perpendicular to this line is one with slope m , such that m*(-2) = -1
so m = 1/2

therefore a general form of the equation that is perpendicular to 4x + 2y = 24 is
y = 0.5x + b
or -x + 2y = 2b
where b is any real number, if we want this perpendicular line to have a common point with 4x + 2y = 24, we can choose (0, 12)
b=12 then since 0 + 2* 12 = 2 * b
one form is -x + 2y = 24

Sincerely Alan